I need help wih my homework, it says " Determine whether each triangle with sides of given lengths is a right triangle. "
The 1st problem is......
1. 6cm, 8cm, 10cm
How do I figure out whether it's right triangle or not ?

Answer by mananth(16946)

If it is a right triangle then Pythagoras theorem will hold good.
6, 8 & 10 are the sides.
the largest side is always the hypotenuse
6^2 +8^2
36+64 = 100
10^2= 100
so 6^2+8^2=10^2
therefore it is a right triangle.